OpenStudy (anonymous):

The volume of a rectangular prism is 115 cm3. If the length is doubled, the width is tripled, and the height is cut in half, what is the volume of the new prism?

OpenStudy (kkutie7):

V=w*h*l=115 V=w^3*h/2*l^2 This might be a good place to start

OpenStudy (anonymous):

width is tripled not cubed, same goes for the length, it's doubled not squared

OpenStudy (anonymous):

V = w*h*l = 115 new V = 3w * h/2 * 2l

OpenStudy (kkutie7):

Shoot you are right my mistake